# Performance Rating Equilibrium

### v2

@@ 3 Applications
In a tournament $T$, there may be multiple fixed points or PREs. However, a PRE can be selected in various ways. In the example below, I numerically calculate the fixed point that converges from an initial vector in which all players have the tournament average rating. While equilibrium selection can also be done using other methods, such as starting with the initial ratings vector instead of the average ratings vector, a PRE obtained with the latter approach has the advantage of not depending on the specific ratings that players had at the start of the tournament. In other words, this approach is entirely performance-based, given the tournament average rating. A numerical approximation for calculating PREs is demonstrated in the following implementation: https://github.com/drmehmetismail/Estimated-Performance-Rating. [^7] FIDE (International Chess Federation) uses certain approximations and specific rules when updating ratings and calculating performance ratings. For a more detailed discussion, see Ismail (2023), who gives a definition of performance ratings for all possible scores.
@@
[…]
Note: Rtg = Rating; Pts = Points; TPR = Tournament Performance Rating; PPR = Perfect Performance Rating.
Table 1 presents the PPRs from two top Swiss tournaments: FIDE Grand Prix Series in Palma de Mallorca and Sharjah. In this table, each player’s PPR is the rating $r_{i}$ that satisfies the equation
$m_{i}=\sum_{\begin{subarray}{c}j=1\end{subarray}}^{9}\frac{1}{1+10^{\frac{r_{i_{j}}-r_{i}}{400}}}$
where $m_{i}$ is the player’s final score in the tournament. This equation ensures that players’ final scores are equal to their total expected scores. If players had entered the tournament with their calculated PPRs as initial ratings, their final scores would be precisely the scores predicted by these ratings. In other words, each player’s initial PPR perfectly predicts the player’s final score.
@@ 6 Uhlmann 14 2636 18 Matulovic 9 2477
Table 2 shows the PPRs from the 1970 Palma de Mallorca Interzonal Tournament. The Interzonal events were part of the World Chess Championship cycle and are similar to the current Chess World Cup. Since the first official FIDE Elo list was not published until 1971, players in this tournament did not have official ratings. However, based on the first published ratings for the players, it is possible to estimate the tournament’s average rating at approximately 2557. The resulting PPRs are presented in the table.

### v3

@@ 4 Applications
There may be multiple PREs due to translation invariance. However, there are various ways in which a PRE can be selected. In the examples below, we numerically calculate the fixed point that converges from an initial vector in which all players have the tournament average rating. [^8] All ratings in the examples are rounded. While equilibrium selection can also be done using other methods, such as starting with the initial ratings vector instead of the average ratings vector, a PRE obtained with the latter approach has the advantage of not depending on the specific ratings that players had at the start of the tournament. [^9] To be sure, a PRE is generally independent of the initial ratings; though this information can be used for equilibrium selection. In other words, this approach is entirely performance-based, given the tournament average rating. A numerical approximation for calculating PREs is demonstrated in the following implementation: https://github.com/drmehmetismail/Perfect-Performance-Ratings. [^10] FIDE (International Chess Federation) uses certain approximations and specific rules when updating ratings and calculating performance ratings. For a more detailed discussion, see Ismail (2023), who gives a definition of performance ratings for all possible scores.
Table 1 presents the PPRs and TPRs for two different three-player round-robin tournaments. Let the 2450-rated player be denoted as A, the 2200-rated player as B, and the 2000-rated player as C. In Tournament 1, the final ranking is the reverse of the initial ranking, indicating that the pre-tournament ratings do not accurately reflect the players’ actual performances. As a result, there is a stark contrast between the TPRs and PPRs. Here, player C, who is underrated, scores 1.5 points in two games, while player A, who is overrated, scores only 0.5 points. According to Equation 2, the expected score of a 2348-rated player (C) against players rated 2217 and 2085 is 1.5. Similarly, the expected score of a 2085-rated player (A) against players rated 2217 and 2348 is 0.5.
In summary, the PPRs are set such that players’ expected scores, based on their opponents’ PPRs, equal their actual points in the tournament. These values are obtained by numerically calculating the fixed point, converged from the initial vector of ratings (2219, 2219, 2219), where 2219 is the tournament’s average rating. [^11] Convergence to the fixed point was achieved after eight iterations.
Tournament 1 Tournament 2
Rating Points TPR PPR Rating Points TPR PPR
2000 1.5 2538 2348 2450 1.5 2305 2348
2200 1.0 2225 2217 2000 1.0 2325 2217
2450 0.5 1895 2085 2200 0.5 1961 2085
Table 1: Perfect Performance Ratings in three-player round-robin tournaments with different outcomes
In Tournament 2, there are also significant differences between the TPRs and PPRs. Notably, player C’s TPR is higher than player A’s TPR, even though C scores fewer points than A. This discrepancy arises in part because player A is overrated; that is, A’s pre-tournament rating is too high relative to A’s final score. For player B, TPR is lower than PPR, as the PPR reflects the strong performance of player C, who is significantly underrated. As this table illustrates, PPRs are entirely based on actual tournament performance and the tournament average rating, without being affected by specific pre-tournament ratings.
@@
Table 2 presents the PPRs from two top Swiss tournaments: FIDE Grand Prix Series in Palma de Mallorca and Sharjah. If players had entered the tournament with their calculated PPRs as initial ratings, their final scores would be equal to the scores predicted by these ratings. In other words, each player’s initial PPR perfectly predicts the player’s final score.
@@ 6 Uhlmann 14 2636 18 Matulovi 9 2477
Table 3 shows the PPRs from the 1970 Palma de Mallorca Interzonal Tournament. The Interzonal events were part of the World Chess Championship cycle and are similar to the current Chess World Cup. Since the first official FIDE Elo list was not published until 1971, players in this tournament did not have official ratings. However, because PPRs do not require pre-tournament ratings, estimating the tournament’s average rating is sufficient for calculating PPRs. We calculated the average rating of the participating players based on their published 1971 ratings and used this estimate. The resulting PPRs are presented in the table. Remarkably, Fischer achieves a PPR of 2805, more than 135 Elo points above the second-ranked player.
5 Concluding remarks
In this note, we introduced a novel method for measuring player performance in tournaments called Perfect Performance Ratings (PPRs). Unlike the Tournament Performance Rating (TPR), PPRs take into account players’ actual performance and achieve perfect consistency between predicted and observed outcomes
While chess served as the primary example in this note, PPRs have potential applications across various domains where Elo-based ratings are commonly used. Future research could apply PPRs to other sports, dating applications, online recommender systems, and evaluations of machine learning models, such as large language models. As Elo (1978) suggested regarding TPR, PPRs can similarly function as a robust rating system that can be calculated across predefined time periods.